OpenStudy (anonymous):

If a = 3 and b = 5, what is the value of a 2 + b 2?

OpenStudy (dls):

\[\Huge (a+b)^2=a^2+b^2+2ab\]

OpenStudy (dls):

isolate required term and substitute a and b

OpenStudy (anonymous):

@DLS Are you kidding us ? The question is to calculate : \[\Large a^2+b^2\] Not : \[\Large (a+b)^2\]

OpenStudy (dls):

-_- can't you write it like this \[\Huge a^2+b^2=(a+b)^2-2ab\]

OpenStudy (dls):

64-30

OpenStudy (anonymous):

9+25

OpenStudy (anonymous):

So we have : \[\Large a=3~~~b=5\] then : \[\Large a^2+b^2=3^2+5^2=9+25=34\]